OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6.

OpenStudy (anonymous):

it is one of these

OpenStudy (anonymous):

i dont know which graph though :(

sam (.sam.):

First try arranging that equation to the form y=mx+c

OpenStudy (anonymous):

i dont understand though

sam (.sam.):

y=mx+c tells you almost everything, but you'll have to get something similar to that first, so -x - 2y = 6 add add x to both sides -2y=x+6 Then, divide -2 both sides, \[y=-\frac{x}{2}-3\] You've got something similar to y=mx+c, c is the y-intercept, which means where the line touches the y-axis. and 'c' here is -3, so, the answer should be C or D

OpenStudy (anonymous):

so they answer is c or d but you dont know which?

sam (.sam.):

Well you'll have to find it yourself :p ,Try letting 'y' equals zero (this is where the line touches the x-axis), \[-\frac{x}{2}-3=0\] \[-\frac{x}{2}=3 \\ \\ x=-6\]

sam (.sam.):

The line touches the x axis at x=-6, so, which answer would it be?

OpenStudy (anonymous):

Sam i dont know i really dont ;(

sam (.sam.):

Come on its so easy, its right infront of you, the line touching the horizontal line at -6

OpenStudy (anonymous):

no it is not touching this -6

sam (.sam.):

Um it is, but you can't see it, lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

damn i was gonna say D too